# Rethinking Text-based Protein Understanding: Retrieval or LLM?

### v3

@@ Abstract
In recent years, protein-text models have gained significant attention for their potential in protein generation and understanding. Current approaches focus on integrating protein-related knowledge into large language models through continued pretraining and multi-modal alignment, enabling simultaneous comprehension of textual descriptions and protein sequences. Through a thorough analysis of existing model architectures and text-based protein understanding benchmarks, we identify significant data leakage issues present in current benchmarks. Moreover, conventional metrics derived from natural language processing fail to accurately assess the model’s performance in this domain. To address these limitations, we reorganize existing datasets and introduce a novel evaluation framework based on biological entities. Motivated by our observation, we propose a retrieval-enhanced method, which significantly outperforms fine-tuned LLMs for protein-to-text generation and shows accuracy and efficiency in training-free scenarios. Our code and data can be seen in https://github.com/IDEA-XL/RAPM.
@@
Juntong Wu1, 2  [^1] Equal contribution. This work was done during Juntong’s internship at IDEA. , Zijing Liu2  [^1] Equal contribution. This work was done during Juntong’s internship at IDEA. , He Cao2,  [^1] Equal contribution. This work was done during Juntong’s internship at IDEA.  [^2] Project Lead., Hao Li1, 2, Bin Feng2, Zishan Shu1, Ke Yu1, Li Yuan 1  [^3] Corresponding author., Yu Li2  [^3] Corresponding author., 1 Peking University, 2 International Digital Economy Academy (IDEA) Correspondence: wujt@stu.pku.edu.cn, yuanli-ece@pku.edu.cn, liyu@idea.edu.cn
@@ 1 Introduction
In recent years, large language models (LLMs) have achieved remarkable success across diverse domains (Brown et al., 2020; Touvron et al., 2023; Wei et al., 2022; Sallam, 2023; Li et al., 2025). To further enhance the ability of LLMs in understanding domain-specific data (e.g., chemistry, biology), researchers have extended LLMs into the multi-modal domain, giving rise to multi-modal large language models (MLLMs) (Liu et al., 2023a; Cao et al., 2023; Maaz et al., 2023). Unlike traditional LLMs, which process single textual modality, MLLMs integrate multiple modalities, such as images, text, and graphs, by aligning them within a unified framework. This is typically accomplished using unimodal encoders for each input type and cross-modal projectors that map different modalities into a shared embedding space (Zhu et al., 2023). As a result, MLLMs enable sophisticated cross-modal reasoning, paving the way for applications like image-text understanding and molecule-function analysis Li et al. (2023); Cao et al. (2024); Liu et al. (2023c).
The advances in MLLMs have led to significant developments in text-based protein understanding (Liu et al., 2024c, 2023b; Zhou et al., 2025; Lv et al., 2024; Liu et al., 2024b). The input of most tasks generally consists of a protein sequence paired with natural language text, while the output represents a functional description. Given that proteins can be represented as amino acid sequences, they are naturally compatible with LLMs and can be processed in two primary ways (Fig. 1a): (1) directly as textual inputs to a language model in a decoder-only or encoder-decoder architechture (Fang et al., 2024; Lv et al., 2024; Luo et al., 2022; Pei et al., 2023), or (2) as an external modality, where specialized encoders first extract high-quality protein representations before alignment with LLMs for downstream tasks (Liu et al., 2024b). To evaluate such protein-text multimodal models, several benchmarks have been introduced, covering key tasks such as protein function prediction, subcellular localization, catalytic activity, and protein design (Fang et al., 2024). The model’s performance is assessed by comparing the model’s outputs against ground-truth annotations. While promising, existing methods raise key questions:
@@
To answer the two questions, we recall that retrieval methods have long served as fundamental approaches in protein tasks, leveraging sequence alignment and database search techniques to identify functional and structural similarities (Lee et al., 2007; Higdon et al., 2010; Eswar et al., 2006). These well-established methods provide a natural baseline for evaluating whether modern LLMs offer genuine advances in protein understanding or merely replicate retrieval paradigms through alternative mechanisms. We therefore tackle Q1 by first comparing traditional retrieval methods against LLMs. Surprisingly, our analysis reveals that simple retrieval-based approaches can match or even outperform current LLMs in protein sequence understanding, challenging the prevailing view that LLMs are inherently superior in this domain.
Through a comprehensive analysis of prevailing protein-text datasets and evaluation metrics, we identify two key limitations in current benchmarks: (1) significant data leakage issues that compromise benchmark validity, and (2) metrics that fail to adequately capture model performance on biologically meaningful tasks. We systematically evaluate both fine-tuning-based and alignment-based approaches across existing datasets, revealing that MLLMs primarily generate outputs by memorizing and reproducing similar input features. Motivated by our analysis and findings, we propose a more rigorous benchmark for text-based protein understanding and introduce an efficient protein knowledge retrieval system, which achieves the state-of-the-art performance in protein understanding by Retrieval-Augmented Protein Modeling (RAPM).
@@ 2.1 Language Models in Protein
Protein language models (PLMs) have successfully adapted Transformer architectures to represent protein sequences as biological tokens, enabling advances in protein embedding Hayes et al. (2025); Brandes et al. (2022); Elnaggar et al. (2021); Cao and Shen (2021); Hu et al. (2024); Chen et al. (2024a, b) and design Madani et al. (2023); Nijkamp et al. (2023); Lv et al. (2024); Ferruz et al. (2022). However, their inability to integrate textual information limits cross-modal reasoning, prompting recent work to develop mixed protein-text models. These approaches include Contrastive Learning Methods Xu et al. (2023); Wu et al. (2024) that align protein sequences with text, Bioknowledge-Augmented Pre-training Ferruz et al. (2022); Taylor et al. (2022); Lv et al. (2024); Pei et al. (2023); Zhuo et al. (2024); Liu et al. (2024b) that leverage large protein-text corpora, and Multi-modal LLMs Liu et al. (2024c); Abdine et al. (2024); Wang et al. (2024); Chen et al. (2024b); Ma et al. (2025); Xiang et al. (2024) that project protein embeddings into LLM spaces. Despite progress, scaling these methods to larger LLMs remains challenging due to prohibitive retraining costs and catastrophic forgetting Kirkpatrick et al. (2017), motivating research into parameter-efficient adaptation strategies.
@@ 2.2 Related Benchmarks
To advance research on protein-text hybrid models, several relevant benchmarks have been proposed. These benchmarks can be categorized into two types: (1) Protein Captioning Tasks, where only the protein sequence is input and a corresponding textual description is generated (e.g., the Swiss-Prot Bairoch and Apweiler (2000) and ProteinKG datasets Zhang et al. (2022)), and (2) Protein Question-Answering Tasks, where both a protein sequence and a question are provided as input, and the model must generate an answer based on the protein and the query (e.g., Mol-Instructions Fang et al. (2024), UniProtQA Luo et al. (2024)), ProteinLMBench Shen et al. (2024). To evaluate model performance on these benchmarks, researchers typically employ standard NLP metrics such as ROUGE Lin (2004), BLEU Papineni et al. (2002), and METEOR Banerjee and Lavie (2005) to measure the similarity between predicted answers and ground-truth references. Models that perform well on these tasks can be applied to automated protein annotation, protein design, and protein property-related QA, thereby facilitating progress in the field.
@@ 2.3 Protein Related Retrieval-Based Methods
Retrieval-based approaches are fundamental to protein science, grounded in the well-established biological principle that sequence homology implies evolutionary conservation and functional similarity Pearson (2013). Single-sequence alignment approaches Altschul et al. (1990); Buchfink et al. (2015); Steinegger and Söding (2017); van Kempen et al. (2022) and multiple sequence alignment tools Remmert et al. (2012); Johnson et al. (2010) are extensively used in bioinformatics for identifying highly homologous sequences. Many protein models utilize retrieval methods to assist downstream tasks, where AlphaFold2 Jumper et al. (2021), MSA-Transformer Rao et al. (2021), and RosettaFold Baek et al. (2021) employ multiple sequence alignment results to aid property prediction or structure folding. Furthermore, retrieval-based approaches Tan et al. (2024); Shaw et al. (2024); Sgarbossa and Bitbol (2025); Jin et al. (2024); Li et al. (2024); Ma et al. (2024) have demonstrated the feasibility of using retrieval tools to enhance LLM-based predictions in protein research.
@@ 3 Analysis
Table 1: Performance comparison of LLM-based and retrieval-based methods across two text-based protein understanding benchmarks. Arch. denotes the model architecture. SFT indicates whether the model has undergone supervised fine-tuning on the training set. Bold for the best. Underline for the second.
@@
For LLM-based methods, we test a variety of model architectures, including 5 unimodal LLMs (Galactica-1.3B-SFT Taylor et al. (2022), BioGPT-347M-SFT Luo et al. (2022), ProLLaMA-7B-SFT Lv et al. (2024), Mol-Instructions-7B Fang et al. (2024), and Llama-3.1-8B-SFT Dubey et al. (2024)), 1 encoder-decoder model (BioT5-Plus Pei et al. (2023)), and 2 multi-modal LLMs (EvoLLAMA-8.8B Liu et al. (2024b) and FAPM-10B Xiang et al. (2024)). For retrieval approaches, we employ MMSeqs2 Steinegger and Söding (2017) for sequence retrieval and ESM-2-650M Lin et al. (2022) as the protein sequence encoder for embedding similarity. Our evaluation results (Table 1) highlight a key finding: all current deep learning methods underperform retrieval-based approaches on these benchmarks. We find that multi-modal LLMs merely match the performance of retrieval methods, while unimodal LLMs demonstrate poorer results. More critically, fine-tuning LLMs requires significant GPU resources, whereas ESM2-based retrieval only needs to compute protein embeddings, and MMSeqs2 retrieval completes 100 million comparisons within 1 minute using only one CPU.
To investigate why the retrieval-based methods beat LLMs, we first examine the data leakage in current benchmarks. The t-SNE visualization of the ESM2 embeddings of the proteins reveals samples forming distinct clusters with significant training-test contamination (Figure 3). We then quantify the level of data leakage by the percentage of test samples whose label can be directly obtained by retrieving the most similar sample (right table in Figure 3). It is easy to see that the leakage rates exceed 50% for most tasks, surpass 95% in some extreme cases. Such pervasive leakage suggests that models fine-tuned on these benchmarks predominantly memorize dataset-specific features rather than develop meaningful biological understanding.
@@ 3.3 Is Retrieval a Silver Bullet?
Table 2: Performance degradation of retrieval methods with full corpus as candidates pool.
@@ 4 Methods: Combine Retrieval & LLM and New Benchmark
To address this need, we develop Retrieval-Augmented Protein Modeling (RAPM) based on the Retrieval-Augmented Generation (RAG) paradigm. RAG is a proven approach for enhancing LLM factual accuracy and domain knowledge Lewis et al. (2020). Our method leverages a Bio-Knowledge Database and contextual prompts to provide LLMs with explicit protein evidence during inference, thereby improving their understanding of biological information and addressing the memorization vs. reasoning tradeoff inherent in this problem.
@@ 4.1 Protein Knowledge Database Construction
For optimal performance, an accurate and efficient domain retrieval system relies on a carefully curated protein knowledge database. In our database, existing biological annotations are standardized into structured [Protein, Annotation] tuples, indexed by amino acid sequence and embeddings. The details of the database construction is shown in Fig. 5, including:
  • Protein Annotation Data Collection. We gather protein-annotation data from InterPro Blum et al. (2025), EC-GO Gligorijević et al. (2021), and Mol-Instructions Fang et al. (2024), extracting biological entity annotations using the method in Sec. 4.3.
@@ 4.2 Retrieval-Augmented Protein Modeling
Formally, given a protein query $\mathcal{Q}$, we retrieve $K$ support data points $\{d_{i}\}_{i=1}^{K}$ from this structured database by ranking candidate entries based on a similarity score $s_{i}$. This score is a weighted combination of sequence and embedding similarity: $s_{i}=\alpha\cdot\mathrm{sim}_{seq}(s,s_{i})+(1-\alpha)\cdot\mathrm{sim}_{emb}(\mathbf{e},\mathbf{e}_{i})$, where $s$ and $s_{i}$ are protein sequences, $\mathbf{e}$ and $\mathbf{e}_{i}$ are the corresponding ESM-2 embeddings, and $\alpha$ is a weight parameter, currently set to $0.5$. Instead of including full sequences in the prompt, each of the top-$K$ retrieved items is formatted as a concise [Confidence, Annotation] tuple. The Confidence level is derived from $s_{i}$ based on quantiles: $>90\%$ as High , $90\%>s_{i}>60\%$ as Medium, and $\leq 60\%$ as Low. The final input prompt $\mathcal{P}$ for the LLM is constructed by concatenating the query, few-shot examples, and the formatted retrieved items: $\mathcal{P}=\mathcal{Q}\oplus\mathcal{E}_{\text{few-shot}}\oplus\mathcal{R}_{1:k}$, where $\mathcal{R}_{1:k}$ represents the formatted top-$K$ entries and $\mathcal{E}_{\text{few-shot}}$ are demonstrations from the training dataset included to help the LLM understand the task format and reasoning. The LLM is then conditioned on $\mathcal{P}$ to predict the answer: $\hat{y}=\mathrm{LLM}(\mathcal{P})$.

### v4

@@ Abstract
In recent years, protein-text models have gained significant attention for their potential in protein generation and understanding. Current approaches focus on integrating protein-related knowledge into large language models through continued pretraining and multi-modal alignment, enabling simultaneous comprehension of textual descriptions and protein sequences. Through a thorough analysis of existing model architectures and text-based protein understanding benchmarks, we identify significant data leakage issues present in current benchmarks. Moreover, conventional metrics derived from natural language processing fail to assess the model’s performance in this domain accurately. To address these limitations, we reorganize existing datasets and introduce a novel evaluation framework based on biological entities. Motivated by our observation, we propose a retrieval-enhanced method, which significantly outperforms fine-tuned LLMs for protein-to-text generation and shows accuracy and efficiency in training-free scenarios. Our code and data can be seen in https://github.com/IDEA-XL/RAPM.
@@
Juntong Wu1, 2  [^1] Equal contribution. , Zijing Liu2  [^1] Equal contribution. , He Cao2  [^1] Equal contribution.  [^2] Project Lead., Hao Li1, 2, Bin Feng2, Zishan Shu1, Ke Yu1  [^3] Corresponding author., Li Yuan 1  [^3] Corresponding author., Yu Li2  [^3] Corresponding author., 1 Peking University, Shenzhen Graduate School 2 International Digital Economy Academy (IDEA) Correspondence: yuke.sz@pku.edu.cn, yuanli-ece@pku.edu.cn, liyu@idea.edu.cn
@@ 1 Introduction
In recent years, large language models (LLMs) have achieved remarkable success across diverse domains (brown2020language; touvron2023llama; wei2022emergent; sallam2023utility; li2025decoupled). To further enhance the ability of LLMs in understanding domain-specific data (e.g., chemistry, biology), researchers have extended LLMs into the multi-modal domain, giving rise to multi-modal large language models (MLLMs) (liu2023visual; cao2023instructmol; maaz2023video). Unlike traditional LLMs, which process single textual modality, MLLMs integrate multiple modalities, such as images, text, and graphs, by aligning them within a unified framework. This is typically accomplished using unimodal encoders for each input type and cross-modal projectors that map different modalities into a shared embedding space (zhu2023minigpt). As a result, MLLMs enable sophisticated cross-modal reasoning, paving the way for applications like image-text understanding and molecule-function analysis li2023blip; cao2024presto; liu2023molca.
The advances in MLLMs have led to significant developments in text-based protein understanding (liu-etal-2024-prott3; proteindt; evolla; lv2024prollama; liu2024evollama). The input of most tasks generally consists of a protein sequence paired with natural language text, while the output represents a functional description. Given that proteins can be represented as amino acid sequences, they are naturally compatible with LLMs and can be processed in two primary ways (Fig. 1a): (1) directly as textual inputs to a language model in a decoder-only or encoder-decoder architecture (fang2023mol; lv2024prollama; biogpt; BioT5), or (2) as an external modality, where specialized encoders first extract high-quality protein representations before alignment with LLMs for downstream tasks (liu2024evollama). To evaluate such protein-text multimodal models, several benchmarks have been introduced, covering key tasks such as protein function prediction, subcellular localization, catalytic activity, and protein design (fang2023mol). The model’s performance is assessed by comparing the model’s outputs against ground-truth annotations. While promising, existing methods raise key questions:
@@
To answer the two questions, we recall that retrieval methods have long served as fundamental approaches in protein tasks, leveraging sequence alignment and database search techniques to identify functional and structural similarities (lee2007predicting; higdon2010modeling; eswar2006comparative). These well-established methods provide a natural baseline for evaluating whether modern LLMs offer genuine advances in protein understanding or merely replicate retrieval paradigms through alternative mechanisms. We therefore tackle Q1 by first comparing traditional retrieval methods against LLMs. Surprisingly, our analysis reveals that simple retrieval-based approaches can match or even outperform current LLMs in protein sequence understanding, challenging the prevailing view that LLMs are inherently superior in this domain.
Through a comprehensive analysis of prevailing protein-text datasets and evaluation metrics, we identify two key limitations in current benchmarks: (1) significant data leakage issues that compromise benchmark validity, and (2) metrics that fail to adequately capture model performance on biologically meaningful tasks. We systematically evaluate both LLM-based and retrieval-based approaches across existing datasets, revealing that MLLMs primarily generate outputs by memorizing and reproducing similar input features. Motivated by our analysis and findings, we propose a more rigorous benchmark for text-based protein understanding and introduce an efficient protein knowledge retrieval system, which achieves the state-of-the-art performance in protein understanding by Retrieval-Augmented Protein Modeling (RAPM).
@@ 2.1 Language Models in Protein
Protein language models (PLMs) have successfully adapted Transformer architectures to represent protein sequences as biological tokens, enabling advances in protein embedding doi:10.1126/science.ads0018; brandes2022proteinbert; elnaggar2021prottrans; cao2021tale; ProtGO; chen2024xtrimopglm; chen2024unifying; xue2022multimodal and design madani2023large; nijkamp2023progen2; lv2024prollama; ferruz2022protgpt2. However, their inability to integrate textual information limits cross-modal reasoning, prompting recent work to develop mixed protein-text models. These approaches include Contrastive Learning Methods protst; wu2024proteinclip that align protein sequences with text, Bioknowledge-Augmented Pre-training ferruz2022protgpt2; taylor2022galactica; lv2024prollama; BioT5; zhuo2024protllm; liu2024evollama that leverage large protein-text corpora, and Multi-modal LLMs liu-etal-2024-prott3; abdine2024prot2text; wang2024protchatgpt; chen2024unifying; ma2025prottex; xiang2024fapm that project protein embeddings into LLM spaces. Despite progress, scaling these methods to larger LLMs remains challenging due to prohibitive retraining costs and catastrophic forgetting kirkpatrick2017overcoming, motivating research into parameter-efficient adaptation strategies.
@@ 2.2 Related Benchmarks
To advance research on protein-text hybrid models, several relevant benchmarks have been proposed. These benchmarks can be categorized into two types: (1) Protein Captioning Tasks, where only the protein sequence is input and a corresponding textual description is generated (e.g., the Swiss-Prot bairoch2000swiss and ProteinKG datasets zhang2022ontoprotein), and (2) Protein Question-Answering Tasks, where both a protein sequence and a question are provided as input, and the model must generate an answer based on the protein and the query (e.g., Mol-Instructions fang2023mol, UniProtQA luo2024biomedgpt, ProteinLMBench shen2024fine, Prot2Text abdine2024prot2text). To evaluate model performance on these benchmarks, researchers typically employ standard NLP metrics such as ROUGE lin2004rouge, BLEU papineni2002bleu, and METEOR banerjee2005meteor to measure the similarity between predicted answers and ground-truth references. Models that perform well on these tasks can be applied to automated protein annotation, protein design, and protein property-related QA, thereby facilitating progress in the field.
@@ 2.3 Protein Related Retrieval-Based Methods
Retrieval-based approaches are fundamental to protein science, grounded in the well-established biological principle that sequence homology implies evolutionary conservation and functional similarity pearson2013introduction. Single-sequence alignment approaches altschul1990basic; buchfink2015fast; steinegger2017mmseqs2; van2022foldseek and multiple sequence alignment tools remmert2012hhblits; johnson2010hidden are extensively used in bioinformatics for identifying highly homologous sequences. Many protein models utilize retrieval methods to assist downstream tasks, where AlphaFold2 jumper2021highly, MSA-Transformer rao2021msa, and RosettaFold baek2021accurate employ multiple sequence alignment results to aid property prediction or structure folding. Furthermore, retrieval-based approaches tan2024retrieval; shaw2024protex; sgarbossa2025rag; jin2024prollm; li2024retrieval; ma-etal-2024-retrieved have demonstrated the feasibility of using retrieval tools to enhance LLM-based predictions in protein research.
@@ 3 Analysis
Table 1: Performance comparison of LLM-based and retrieval-based methods across two text-based protein understanding benchmarks. Arch. denotes the model architecture. SFT indicates whether the model has undergone supervised fine-tuning on the training set. Bold denotes the best. Underline denotes the second best.
@@
For LLM-based methods, we test a variety of model architectures, including 5 unimodal LLMs (Galactica-1.3B-SFT taylor2022galactica, BioGPT-347M-SFT biogpt, ProLLaMA-7B-SFT lv2024prollama, Mol-Instructions-7B fang2023mol, and Llama-3.1-8B-SFT llama3), 1 encoder-decoder model (BioT5-Plus BioT5), and 2 multi-modal LLMs (EvoLLAMA-8.8B liu2024evollama and FAPM-10B xiang2024fapm). For retrieval approaches, we employ MMSeqs2 steinegger2017mmseqs2 for sequence retrieval and ESM-2-650M esm2 as the protein sequence encoder for embedding similarity. Our evaluation results (Table 1) highlight a key finding: all current deep learning methods underperform retrieval-based approaches on these benchmarks. We find that multi-modal LLMs merely match the performance of retrieval methods, while unimodal LLMs demonstrate poorer results. More critically, fine-tuning LLMs requires significant GPU resources, whereas ESM2-based retrieval only needs to compute protein embeddings, and MMSeqs2 retrieval completes 100 million comparisons within 1 minute using only one CPU.
To investigate why the retrieval-based methods beat LLMs, we first examine the data distribution in current benchmarks. The t-SNE visualization of the ESM2 embeddings of the proteins reveals samples forming distinct clusters with significant training-test contamination (Figure 3). Prot2Text is the multimodal dataset collected by abdine2024prot2text, where they reduced the overlap between the training and test samples. We then quantify the level of label leakage by the percentage of test samples whose label can be directly obtained by retrieving the most similar sample (right table in Figure 3). It is easy to see that the leakage rates exceed 50% for most tasks, surpassing 95% in some extreme cases. The process of protein function annotation possibly causes such pervasive label leakage and suggests that models fine-tuned on these benchmarks predominantly memorize dataset-specific features rather than develop meaningful biological understanding.
@@ 3.3 Is Retrieval a Silver Bullet?
Table 2: Performance degradation of retrieval methods with the full corpus as the candidate pool.
@@ 4 Methods: Combine Retrieval & LLM and New Benchmark
To address this need, we develop Retrieval-Augmented Protein Modeling (RAPM) based on the Retrieval-Augmented Generation (RAG) paradigm. RAG is a proven approach for enhancing LLM factual accuracy and domain knowledge lewis2020retrieval. Our method leverages a Bio-Knowledge Database and contextual prompts to provide LLMs with explicit protein evidence during inference, thereby improving their understanding of biological information and addressing the memorization vs. reasoning tradeoff inherent in this problem.
@@ 4.1 Protein Knowledge Database Construction
For optimal performance, an accurate and efficient domain retrieval system relies on a carefully curated protein knowledge database. In our database, existing biological annotations are standardized into structured [Protein, Annotation] tuples, indexed by amino acid sequence and embeddings. The details of the database construction are shown in Fig. 5, including:
  • Protein Annotation Data Collection. We gather protein-annotation data from InterPro InterPro, EC-GO gligorijevic2021structure, and Mol-Instructions fang2023mol, extracting biological entity annotations using the method in Sec. 4.3.
@@ 4.2 Retrieval-Augmented Protein Modeling
Formally, given a protein query $\mathcal{Q}$, we retrieve $K$ support data points $\{d_{i}\}_{i=1}^{K}$ from this structured database by ranking candidate entries based on a similarity score $s_{i}$. This score is a weighted combination of sequence and embedding similarity: $s_{i}=\alpha\cdot\mathrm{sim}_{seq}(s,s_{i})+(1-\alpha)\cdot\mathrm{sim}_{emb}(\mathbf{e},\mathbf{e}_{i})$, where $s$ and $s_{i}$ are protein sequences, $\mathbf{e}$ and $\mathbf{e}_{i}$ are the corresponding ESM-2 embeddings, and $\alpha$ is a weight parameter, currently set to $0.5$. Instead of including full sequences in the prompt, each of the top-$K$ retrieved items is formatted as a concise [Confidence, Annotation] tuple. The Confidence level is derived from $s_{i}$ based on quantiles: $90\%$ as High , $90\%>s_{i}>60\%$ as Medium, and $\leq 60\%$ as Low. The final input prompt $\mathcal{P}$ for the LLM is constructed by concatenating the query, few-shot examples, and the formatted retrieved items: $\mathcal{P}=\mathcal{Q}\oplus\mathcal{E}_{\text{few-shot}}\oplus\mathcal{R}_{1:k}$, where $\mathcal{R}_{1:k}$ represents the formatted top-$K$ entries and $\mathcal{E}_{\text{few-shot}}$ are demonstrations from the training dataset included to help the LLM understand the task format and reasoning. The LLM is then conditioned on $\mathcal{P}$ to predict the answer: $\hat{y}=\mathrm{LLM}(\mathcal{P})$.